# Simulated Interactive Debugging

### v1

@@
Abstract.
Debugging software, i.e., the localization of faults and their repair, is a main activity in software engineering. Therefore, effective and efficient debugging is one of the core skills a software engineer must develop. However, the teaching of debugging techniques is usually very limited or only taught in indirect ways, e.g., during software projects. As a result, most Computer Science (CS) students learn debugging only in an ad-hoc and unstructured way. In this work, we present our approach called Simulated Interactive Debugging that interactively guides students along the debugging process. The guidance aims to empower the students to repair their solutions and have a proper ”learning” experience. We envision that such guided debugging techniques can be integrated into programming courses early in the CS education curriculum. To perform an initial evaluation, we developed a prototypical implementation using traditional fault localization techniques and large language models. Students can use features like the automated setting of breakpoints or an interactive chatbot. We designed and executed a controlled experiment that included this IDE-integrated tooling with eight undergraduate CS students. Based on the responses, we conclude that the participants liked the systematic guidance by the assisted debugger. In particular, they rated the automated setting of breakpoints as the most effective, followed by the interactive debugging and chatting, and the explanations for how breakpoints were set. In our future work, we will improve our concept and implementation, add new features, and perform more intensive user studies.
@@ 1. Introduction
Despite the paradigm shift towards Artificial Intelligence (AI)-assisted software development, particularly with the emergence of Large Language Models (LLMs), there remains a high demand for qualified software engineers capable of developing reliable, high-quality software. With more auto-generated code, it will become even more critical that junior software developers can effectively debug software and solve bugs, even in code they have not written themselves. Radermacher et al. (Radermacher et al., 2014) explored knowledge deficiencies of graduate students from an industry perspective. One of the most frequently mentioned issues with software tools has been version control systems and debuggers. While we have seen the integration of version control systems like Git in the Computer Science (CS) and Software Engineering (SE) curriculum, e.g., with the usage of GitHub Classroom, we have not seen much innovation regarding debugging education. Michaeli and Romeike (Michaeli and Romeike, 2019) noted that there are only few studies investigating the ”explicit teaching of debugging” and students are often left alone to learn debugging ”the hard way”. Our experience confirms this observation: most students still learn (interactive) debugging of code in an ad-hoc and unstructured way by trial and error. Using ”print/log” statements is often the only extent of students’ experience in debugging. They do not receive proper guidance in efficient program comprehension and how to debug programs effectively. Current CS/SE research efforts focus on general programming education, which helps to provide scalable alternatives to mentor and guide a rising number of students. Automated Program Repair (APR) techniques (Gulwani et al., 2018; Hu et al., 2019; Zhang et al., 2024) can help to produce patches for incorrect submissions, which then can be used as a basis for personalized feedback, or even automated grading (Fan et al., 2023). In fact, we already have seen the application of such techniques in the CS/SE curriculum focusing on repairing students’ solutions and providing feedback (Kazemitabaar et al., 2024b; Fan et al., 2024). However, the existing related work does not address actual debugging.
Our research objective is to understand students’ needs in learning to debug and provide automated means to guide them effectively along the debugging process.
As part of this objective, we strive to seek answers to questions like ”How can we guide the debugging process using artifacts generated from APR/SE techniques?” and ”How can LLMs be integrated into the debugging process for the benefit of the student?”
@@
As a first step to achieve our research objective, we propose the concept of Simulated Interactive Debugging. The essence of this concept is to create a simulation of an interactive debugging experience; ”simulated” because the solution is known in the education context due to given reference implementations and artifacts like fault locations and patches that can be auto-generated with SE and APR techniques. Our goal is to provide a controlled but supportive environment, in which the students debug the problem on their own. Similar to a human tutor who would not directly reveal the solution and instead help the student to understand and fix the issue in their code. Our vision of such an assisted debugger requires integration into coding tools like Integrated Development Environments (IDEs) to guide students actively during the debugging process. Based on APR, fault localization results, and the usage of LLMs, we can show potential program problematic locations, explain the problem at these locations, set breakpoints automatically, identify relevant variable values, highlight them, and provide partial or complete hints on how to repair the student’s program.
@@
As part of a design science methodology (Hevner et al., 2004; Wieringa, 2014), we developed a prototype of the proposed concept as a Visual Studio Code IDE extension. We conducted a pilot user study with eight undergraduate CS students. The results show that students have little experience with (interactive) debugging but enjoy the provided guidance, notably the automatic breakpoint setting and the interaction with the AI-based chatbot. In summary, our core contributions are:
@@ 2. Related Work
Due to the rise in the number of CS students (Masters, 2011), educators need help to cope with the number of assignment submissions and the needed feedback. APR techniques are currently being explored to supplement the missing human tutor, e.g., by realizing so-called Intelligent Tutoring Systems (ITS). Early works in this domain attempt to repair student submissions to generate feedback (Gulwani et al., 2018; Hu et al., 2019) and automatically grade programming assignments (Fan et al., 2023; Birillo et al., 2022). More recently, LLM-based approaches have been explored intensively to generate personalized feedback and hints for programming assignments (Zhang et al., 2024; Liffiton et al., 2024; Hou et al., 2024; Zhao et al., 2024; S Kumar et al., 2024). Several experiences have been reported where LLM-based programming assistants have been deployed in the classroom context (Abolnejadian et al., 2024; Kazemitabaar et al., 2024b; Fan et al., 2024; Kurniawan et al., 2023). LLMs also can be combined with traditional techniques like static analysis to provide incremental hints for programming tasks (Birillo et al., 2024a). Moreover, existing works (Koutcheme et al., 2024; Phung et al., 2023) benchmarked various models regarding their capabilities to support programming education. Overall, the existing works focus on repairing programming assignments and generating student feedback; they do not address the teaching of interactive debugging.
@@
To get an overview of the teaching of debugging, we refer to McCauley et al. (McCauley et al., 2008), who conducted a systematic literature review. Michaeli and Romeike (Michaeli and Romeike, 2019) recently explored the influence of teaching systematic debugging concepts with an intervention study concluding that explicitly teaching debugging skills positively affects debugging self-efficacy. Recent LLM-inspired, debugging-related research focuses on Socratic questioning (Al-Hossami et al., 2024) and generating so-called debugging quizzes (Phung et al., 2023; Padurean et al., 2024) that ask students to reason about a given buggy program and design bug-revealing test cases. While the reported experiences and the existing approaches and tools can help students build general debugging skills, to the best of our knowledge, no approach attempts to guide students through the interactive debugging process.
@@
At the latest, with the emergence of LLMs, AI also arrived in the domain of CS education. Shein (Shein, 2024) argues that while we still need to teach students the fundamentals of programming, such teaching will become more ”alive” through AI, e.g., via chatbots. Denny et al. (Denny et al., 2024) investigated desirable characteristics for AI teaching assistants and concluded that students enjoy AI-based learning support and actually preferred tailored scaffolding instead of mere responses and the revealing of solutions. Kazemitabaar et al. (Kazemitabaar et al., 2024a) explored the design space for such AI tools and highlighted the importance of step-by-step guidance and an interactive dialog with the AI. Following the insights of these previous works, our proposed concept for simulated interactive debugging guides the student in understanding the problem and identifying the programming faults instead of just providing feedback on how to fix it. Further, we also offer an AI-based chatbot with which the students can engage in a controlled environment.
3. Concept: Simulated Interactive Debugging
To tackle the shortcomings of CS education in program debugging, we propose the following concept. Using technology from automated software engineering (SE), in particular from automated program repair (APR), we can identify the issue in a student’s buggy program and can generate a corresponding solution. Note that such a solution can structurally and syntactically differ from the reference implementation provided by a lecturer. Now, knowing the fault and where to fix it, we can use this knowledge to guide the student not only in fixing the issue, but step-by-step in understanding the program’s faults, identifying the potential fix locations, and eventually fixing the problem—similar to how a human tutor would guide the student along the debugging process. Therefore, we designed and developed an AI-assisted debugger that offers students a simulated interactive debugging experience. Students receive help in running tests, using the interactive debugger in identifying the underlying fault in the problem, and in fixing the actual problem.
@@ 3. Concept: Simulated Interactive Debugging
This concept fits into a larger research endeavor of ours, where we aim to develop an active learning environment integrating APR through LLMs and other SE tools with interactive debugging. This concept is designed to provide interactive guidance to students, thereby enhancing their learning experience in programming. Furthermore, our concept exposes students to AI tools, in particular LLMs, in a guided and controlled way. LLMs can be deployed in various roles in this concept, e.g., to generate hints and explanations in natural language, as well as core technology for an interactive chatbot. Based on the generated artifacts from techniques like spectrum-based fault localization (SBFL) (Wong et al., 2016), static code analysis, symbolic execution, code synthesis and LLM prompting, we can set breakpoints automatically, identify relevant variable values, and highlight them for the students. Such a process can also be supported by Socratic questioning (Al-Hossami et al., 2024).
@@
The concept has two types of users: the students and the lecturers, although the main focus is on the students. Both user groups provide inputs and receive feedback. The lecturer provides the programming assignment, the reference implementation, and a test suite consisting of private and public test cases. The student provides a buggy solution and (optionally) a set of custom test cases. There are three different use cases: (1) the primary use case is the assisted debugging (discussed in Section 3.1), (2) the test assessment for lecturers, and (3) the test assessment for custom tests from the student. The test assessment for lecturers checks whether the given test suite is strong enough to catch errors and perform effective fault localization. This can be realized via mutation testing, and additional tests could be generated, e.g., via evolutionary test generation (Fraser and Arcuri, 2011). By offering additional features for the lecturer, our goal is also to increase confidence in the overall system, which is crucial for its successful deployment in programming courses. The test assessment for customs tests from the student works with the idea that when the student submits custom tests, these can be executed against the reference implementation from the lecturer. If they fail on the reference implementation, they are flagged and can also be corrected (because the correct output for the inputs of such test cases could be extracted via the reference implementation). To help students, we also can propose new test cases to them, e.g., to explore certain boundary cases or to increase overall code coverage.
Deployment
Birillo et al. (Birillo et al., 2024b) already argued that combining programming education with the usage of IDEs helps to familiarize students with industrial technologies. Building on top of this idea, our concept is envisioned to be developed as an IDE plugin.
3.1. Assisted Debugging
Figure 1. Envisioned dialogue between student and AI bot.
@@ 3.1. Assisted Debugging
Figure 1 showcases the envisioned interaction between the student and the AI assistant. First, the AI assistant can (1) clarify/explain failing test cases and the observed failures. Next, it can (2) automatically set breakpoints at locations of interest. These locations can be identified, e.g., by using SBFL. Setting such breakpoints enables the unexperienced student to use the interactive debugger to step through the test execution. Based on LLMs, we also can (3) generate hints in natural language explaining why these breakpoints are of interest. In cases where a student’s understanding of the problem is far from what is expected, the AI assistant can (4) deliver incremental guidance, e.g., via progressive hints, to ensure that the student understands the problem correctly. If the student is still stuck, the AI assistant can (5) partially reveal why a certain test case execution fails. Furthermore, we can (6) help the student in the interactive debugging by identifying and highlighting interesting program variables that the student can observe during the debugging session. This interaction aims to guide the student to repair their own solution and to have a proper ”learning” experience.
Figure 2. Screenshot from our prototypical implementation. The left part shows the code and the automatically set breakpoints. The right part shows the feature selection and the dialogue with the AI-based chatbot.
Figure 3. Participants’ programming/debugging experience.
Figure 4. Participants’ assessment of the provided features.
4. Pilot User Study
To provide first insights towards our research goals, we executed a pilot user study. We investigate the practicality of our assisted debugging concept, as well as the usability of our implementation.
4.1. Implementation
@@
As a study prototype, we realized the Simulated Interactive Debugging concept as a Visual Studio Code IDE extension. Our concept is language-agnostic; however, due to the programming courses at our institution, we decided to first focus on supporting Python. It provides two core features for the assisted debugger: the automatic setting of breakpoints (based on SBFL with FauxPy (Rezaalipour and Furia, 2024)) and a chat interface providing tailored debugging hints. Each feature can be triggered via a VSCode command. The chat interface connects the participant with an LLM-powered chatbot. The user can choose from two modes: the Generate Hints mode, which only allows the student to get step-by-step hints, and the Interactive Debugging Guidance mode, which allows the student to engage in a dialogue with the LLM in a controlled way. The student can discuss code issues and query about how to proceed with debugging or improving the code. Figure 2 shows an annotated screenshot of our implementation. As underlying LLM, we used OpenAI gpt-3.5-turbo. Note that in the implementation of our current prototype, we omitted the support for the lecturer’s and the student’s test assessment because our primary interest was in the feedback for the actual debugging.
4.2. Study Setup
@@
The study started with a short briefing, the IRB signing, and the completion of a profiling survey. Then, we spent 15 minutes for providing an introduction to the topic of debugging and for demonstrating the features of our tool. Afterwards, the participants had 40 minutes to solve two coding tasks. The tasks originate from LeetCode and represent programming tasks of medium difficulty that should not overwhelm or underwhelm the students. Both were provided with pre-existing buggy solutions, along with corresponding passing and failing test cases. In the first task, named maxOperations (lee, 2025b), the participants have to fix an algorithm that is supposed to calculate the maximum number of times, to pick two numbers from a given array whose sum equals the parameter $k$. In the second task, named longestOnes (lee, 2025a), the participants have to fix the algorithm that is supposed to compute the maximum number of consecutive 1’s in a binary array if one can flip at most $k$ 0’s. The participants had first to understand the task’s description, then find the errors in the provided solution, and finally provide a modified implementation that passes all test cases. To remember the available assisted debugging features, we also provided each participant with a handout describing a possible workflow to solve the tasks (see replication artifact [^1] Please contact the corresponding author to request the replication artifact.). We informed the participants that the chatbot’s hints might not be entirely accurate. In the end, the participants were asked to complete a usability questionnaire (Brooke, 1996) and a post-task survey. The study was timed for 70 minutes, and the participants were compensated with 30 SGD.
@@
For our pilot study, we recruited eight first-year undergraduate CS students. Based on our profiling survey results, all of them are beginners in the field of programming, and hence, are in our target group. Figure 3 illustrates the self-classified experience levels of our participants: Most of them have programming experience of less than two years and can only write basic Python programs. 6/8 participants mentioned that they can debug basic errors but struggle with complex ones. At the same time, 6/8 have never heard of or never used an interactive debugger. Those who understand the concept of an interactive debugger still prefer using print/log statements to debug their code.
4.3. Results & Discussion
In the following, we discuss the findings from our user study regarding the feedback on the existing features of the assisted debugger, the overall usability of our prototype, the participant’s confidence of manually setting breakpoints, and additionally desired features.

### v2

@@
Abstract
Debugging software, i.e., the localization of faults and their repair, is a key activity in software engineering. Therefore, effective and efficient debugging is one of the core skills a software engineer must develop. However, the teaching of debugging techniques is usually very limited or only taught in indirect ways, e.g., during software projects. As a result, most Computer Science (CS) students learn debugging only in an ad-hoc and unstructured way. In this work, we present our approach called Simulated Interactive Debugging that interactively guides students along the debugging process. The guidance aims to empower the students to repair their solutions and have a proper learning experience. We envision that such guided debugging techniques can be integrated into programming courses early in the CS education curriculum. We developed a prototypical implementation using traditional fault localization techniques and large language models. Students can use features like the automated setting of breakpoints or an interactive chatbot. We designed and executed a small-scale, controlled experiment with eight undergraduate CS students. Based on the responses, we conclude that the participants liked the systematic guidance. They rated the automated setting of breakpoints as most effective, followed by the interactive debugging and chatting, and the breakpoint explanations. In future, we will extend our concept and implementation, and perform more intensive user studies.
Index Terms:
@@ I Introduction
Despite the paradigm shift towards Artificial Intelligence (AI)-assisted software development, particularly with the emergence of Large Language Models (LLMs), there remains a high demand for qualified software engineers capable of developing reliable, high-quality software. With more auto-generated code, it will become even more critical that junior software developers can effectively debug software and solve bugs, even in code they have not written themselves. Radermacher et al. [1] explored knowledge deficiencies of graduate students from an industry perspective. One of the most frequently mentioned issues with software tools has been version control systems and debuggers. While we have seen the integration of version control systems like Git in the Computer Science (CS) and Software Engineering (SE) curriculum, e.g., with the usage of GitHub Classroom, we have not seen much innovation regarding debugging education. Michaeli and Romeike [2] noted that there are only a few studies investigating the “explicit teaching of debugging” and students are often left alone to learn debugging “the hard way”. Our experience confirms this observation: most students still learn (interactive) debugging of code in an ad-hoc and unstructured way by trial and error. Using “print/log” statements is often the only extent of students’ experience in debugging. They do not receive proper guidance in efficient program comprehension and how to debug programs effectively. Current CS/SE research efforts focus on general programming education, which helps to provide scalable alternatives to mentor and guide a rising number of students. Automated Program Repair (APR) techniques [3, 4, 5] can help to produce patches for incorrect submissions, which then can be used as a basis for personalized feedback, or even automated grading [6]. In fact, we already have seen the application of such techniques in the CS/SE curriculum focusing on repairing students’ solutions and providing feedback [7, 8]. However, the existing related work in CS education does not address actual debugging. Therefore, our research objective is to understand students’ needs in learning to debug and provide automated means to guide them effectively along the debugging process.
@@
As a first step to achieve our research objective, we propose Simulated Interactive Debugging. The essence of this concept is to create a simulation of an interactive debugging experience; “simulated” because the solution is known in the education context due to given reference implementations and auto-generated artifacts like fault locations and patches. Our goal is to provide a controlled but supportive environment, in which the students debug the problem on their own. Similar to a human tutor, we would not directly reveal the solution and instead help the student to understand and fix the issue. Our vision of such an assisted debugger requires integration into coding tools like Integrated Development Environments (IDEs) to actively guide students. Based on APR, fault localization, and LLMs, we can show potential problematic locations, explain the issues at these locations, set breakpoints, identify relevant variable values, and provide repair hints.
@@
As part of a design science methodology [9, 10], we developed a prototype as a VS Code IDE extension and conducted a pilot user study with eight undergraduate CS students. The results indicate that students have little experience with (interactive) debugging but enjoy the guidance, notably the automatic breakpoint setting and the interaction with the chatbot. In summary, our core contributions are:
@@ II Related Work
Due to the rise in the number of CS students [11], educators need help to cope with the number of assignment submissions. APR techniques are currently being explored to supplement the missing human tutor, e.g., by realizing so-called Intelligent Tutoring Systems (ITS). Early works in this domain attempt to repair student submissions to generate feedback [3, 4] and automatically grade programming assignments [6, 12]. More recently, LLM-based approaches have been explored to generate personalized feedback and hints for programming assignments [5, 13, 14, 15, 16]. Several experiences have been reported where LLM-based programming assistants have been deployed in the classroom context [7, 8, 17, 18]. LLMs also can be combined with traditional techniques like static analysis to provide incremental hints for programming tasks [19]. Moreover, existing works [20, 21] benchmarked various models regarding their capabilities to support programming education. Overall, the existing works focus on repairing programming assignments and generating feedback; they do not address the teaching of interactive debugging.
@@
To get an overview of the teaching of debugging, we refer to McCauley et al. [22], who conducted a systematic literature review. Michaeli and Romeike [2] recently explored the influence of teaching systematic debugging concepts with an intervention study concluding that explicitly teaching debugging skills positively affects debugging self-efficacy. Recent LLM-inspired, debugging-related research focuses on Socratic questioning [23] and generating so-called debugging quizzes [21, 24] that ask students to reason about a given buggy program and design bug-revealing test cases. While the reported approaches and tools can help students build general debugging skills, to the best of our knowledge, no approach guides through the interactive debugging process.
@@
Most recently, with the emergence of LLMs, AI has arrived in the domain of CS education. Shein [25] argues that while we still need to teach students the fundamentals of programming, such teaching will become more “alive” through AI, e.g., via chatbots. Denny et al. [26] investigated desirable characteristics for AI teaching assistants and concluded that students enjoy AI-based learning support and actually preferred tailored scaffolding instead of mere responses and the revealing of solutions. Kazemitabaar et al. [27] explored the design space for such AI tools and highlighted the importance of step-by-step guidance and an interactive dialog with the AI. Following the insights of these works, our proposed concept for Simulated Interactive Debugging guides the student in understanding the problem and identifying the programming faults instead of just providing feedback on how to fix it. Further, we also offer an AI-based chatbot with which the students can engage in a controlled environment.
III Concept: Simulated Interactive Debugging
To tackle the shortcomings of CS education in program debugging, we propose the following concept. Using technology from automated SE, in particular from APR, we can identify the issue in a student’s buggy program and can generate a corresponding solution. Note that such a solution can structurally and syntactically differ from the reference implementation provided by a lecturer. Now, knowing the fault and where to fix it, we can use this knowledge to guide the student not only in fixing the issue, but step-by-step in understanding the defect, identifying the potential fix locations, and eventually fixing the problem—similar to how a human tutor would guide the student along the debugging process. We designed and developed an AI-assisted debugger that offers students a simulated interactive debugging experience. Students receive help in running tests, using the interactive debugger in identifying the underlying fault in the problem, and in fixing the actual problem. As Birillo et al. [28] have argued, combining programming education with the use of IDEs helps familiarize students with industrial technologies; hence we envision our concept being deployed within IDEs.
@@ III Concept: Simulated Interactive Debugging
This concept fits into a larger research endeavor of ours, where we aim to develop an active learning environment integrating APR through LLMs and other SE tools with interactive debugging. This concept is designed to provide interactive guidance to students, thereby enhancing their learning experience in programming. Furthermore, our concept exposes students to AI tools, in particular LLMs, in a guided and controlled way. LLMs can be deployed in various roles in this concept, e.g., to generate hints and explanations in natural language, as well as core technology for an interactive chatbot. Based on the generated artifacts from techniques like spectrum-based fault localization (SBFL) [29], static code analysis, symbolic execution, code synthesis and LLM prompting, we can set breakpoints automatically, identify relevant variable values, and highlight them for the students. Such a process can also be supported by Socratic questioning [23].
@@
The concept has two types of users: the students and the lecturers, although the main focus is on the students. Both user groups provide inputs and receive feedback. The lecturer provides the programming assignment, the reference implementation, and a test suite consisting of private and public test cases. The student provides a buggy solution and (optionally) a set of custom test cases. There are three different use cases: (1) the primary use case is the assisted debugging, (2) the test assessment for lecturers, and (3) the test assessment for custom tests from the student. The test assessment for lecturers checks whether the given test suite is strong enough to catch errors and perform effective fault localization, e.g., via mutation testing. The test assessment for custom tests from the student works with the idea that when the student submits custom tests, these can be executed against the reference implementation from the lecturer. Failing test cases can be flagged and corrected; further, we can propose new tests, e.g., to explore boundary cases or to increase coverage.
Assisted Debugging
Figure 1: Envisioned conversation between student and the AI assistant.
@@
Figure 1 shows the envisioned interaction between student and AI assistant. First, the AI assistant can (1) explain failing tests and the observed failures. Next, it can (2) automatically set breakpoints at locations of interest. These locations can be identified, e.g., by using SBFL. Setting such breakpoints enables the unexperienced student to use the interactive debugger to step through the test execution. Based on LLMs, we also can (3) generate hints in natural language explaining these breakpoints. In cases where a student’s understanding of the problem is far from what is expected, the AI assistant can (4) deliver incremental guidance, e.g., via progressive hints. If the student is still stuck, the AI assistant can (5) partially reveal why a certain test fails. Furthermore, we can (6) help the student by highlighting interesting variables that the student can observe during debugging. This interaction aims to guide the student to repair their solution and to have a proper learning experience.
IV Pilot User Study
IV-A User Study Setup
Prototype
@@ IV-A User Study Setup
We realize our study prototype as a VS Code IDE extension. Our concept is language-agnostic; however, due to the programming courses at our institution, we decided to first focus on supporting Python. It provides two core features for the assisted debugger: the automatic setting of breakpoints (based on SBFL with FauxPy [30]) and a chat interface providing tailored debugging hints. The chat interface connects the participant with an LLM-powered chatbot using OpenAI gpt-3.5-turbo. Note that in the implementation of our current prototype, we focused only on the debugging use case.
Figure 2: Participants’ assessment of the provided features.
@@
The study started with a short briefing, the IRB signing, and the completion of a profiling survey. Then, we provided a 15-minute introduction into debugging and for demonstrating the tool’s features. Afterwards, the participants had 40 minutes to solve two coding tasks. The tasks originate from LeetCode and are of medium difficulty that should not overwhelm or underwhelm the students. Both were provided with pre-existing buggy solutions, along with corresponding passing and failing tests. In the first task, named maxOperations [31], the participants have to fix an algorithm that is supposed to calculate the maximum number of times to pick two numbers from a given array whose sum equals the parameter $k$. In the second task, named longestOnes [32], the participants have to fix the algorithm that is supposed to compute the maximum number of consecutive 1’s in a binary array if one can flip at most $k$ 0’s. The participants first had to understand the task’s description, find the errors in the provided solution, and finally modify the program to pass all tests. We provided a handout with a possible workflow to solve the tasks with the available assisted debugging features (see artifact). We informed the participants that the generated hints might not be entirely accurate. In the end, the participants were asked to complete a usability questionnaire [33] and a post-task survey. The study was timed for 70 minutes, and the participants were compensated with 30 SGD.
@@
For our pilot study, we recruited eight first-year undergraduate CS students. Based on our profiling survey results, all of them are beginners in the field of programming, and hence, are in our target group: most of them have programming experience of less than two years and can only write basic Python programs. 6/8 participants mentioned that they can debug basic errors but struggle with complex ones. At the same time, 6/8 have never heard of or never used an interactive debugger. Those who understand the concept of interactive debuggers still prefer using print/log statements.
IV-B Results & Discussion